Two point charges of the same magnitude but opposi
Answer #1

To solve this we need to know the equation of electric field \small E=\frac{kq}{r^2} and that the isosceles triangle has two equal sides and two equal angles.

At point M the electric field has only one component in the x-direction.

\small E_M=\frac{k2q}{r^2} , if we name "a" the distance between the charges and P then \small r=acos\alpha

\small E_M=\frac{k2q}{\left (acos\alpha \right )^2}

At point P, the electric field has one component as well in the x-direction because components in y-directions are equal but opposite signs so

\small E_P=\frac{k2q}{ a^2}cos\alpha

Knowing that \small \frac{E_M}{E_P}=7.1 then \small \frac{\frac{k2q}{\left (acos\alpha \right )^2}}{\frac{k2qcos\alpha}{ a^2}}=7.1 so \small \frac{1}{\left (cos\alpha \right )^3}=7.1

\small \sqrt[3]{\frac{1}{7.1}}=cos\alpha and \small \alpha =cos^{-1}\left (\sqrt[3]{\frac{1}{7.1}} \right )=58.65^{\circ}